OpenStudy (anonymous):

A specimen of an alien substance was recently collected on Pandora. When 0.6 kg of water at 30oC was mixed with .8 kg of the alien substance at 60oC, the resulting temperature was 55oC. Given that the specific heat of water is 4180 J/kg oC, calculate the specific heat of this unknown alien substance.

OpenStudy (shamim):

So alien has higher temperature than water. Right?

OpenStudy (shamim):

So alien will leave heat nd water will take heat

OpenStudy (shamim):

And amount of leaving heat of alien will b equal to amount of taking heat by water

OpenStudy (shamim):

Heat taking by water is=m*s*temperature difference

OpenStudy (shamim):

Q1=0.6*4180*(55-30)

OpenStudy (shamim):

Now leaving heat by alien is Q2=m*s*difference of temperature

OpenStudy (shamim):

Q2=8*s*(60-55)

OpenStudy (shamim):

Now Q1=Q2 S=?